# DiGAN: Diffusion-Guided Attention Network for Early Alzheimer’s Disease Detection

### v1

@@ Abstract
Early diagnosis of Alzheimer’s disease (AD) remains a major challenge due to the subtle and temporally irregular progression of structural brain changes in the prodromal stages. Existing deep learning approaches require large longitudinal datasets and often fail to model the temporal continuity and modality irregularities inherent in real-world clinical data. To address these limitations, we propose the Diffusion-Guided Attention Network (DiGAN), which integrates latent diffusion modelling with attention-guided convolutional network. The diffusion model synthesizes realistic longitudinal neuroimaging trajectories from limited training data, enriching temporal context and improving robustness to unevenly spaced visits. The attention-convolutional layer then captures discriminative structural–temporal patterns that distinguish cognitively normal subjects from those with mild cognitive impairment and subjective cognitive decline. Experiments on synthetic and ADNI datasets demonstrate that DiGAN outperforms existing state-of-the-art baselines, showing its potential for early-stage AD detection.
@@ Introduction
  • We demonstrate the effectiveness of DiGAN compared to existing state-of-the-art methods on both the synthetic and ADNI datasets.
@@ 3. Batch Normalization and Output:
Synthetic dataset.
We use a synthetic longitudinal dataset derived from the DELCODE study (Jessen et al. 2018), to simulate realistic neuroimaging trajectories observed in early-stage AD (Table 1). These profiles were generated by modeling intrasubject trajectories using a $2^{nd}$ polynomial function fitted to real distributions, followed by the addition of low-variance Gaussian noise to simulate intersubject variability and measurement uncertainty. Each profile comprises volumetric measures of cortical and subcortical regions, including the entorhinal cortex, parahippocampal gyrus, precuneus, posterior cingulate, and temporal lobes, as well as white matter hyperintensities (WMH) in the frontal and parietal lobes, perivascular spaces (PVS) in the basal ganglia and centrum semiovale.
@@
A longitudinal study widely used for Alzheimer’s disease research. Subjects with at least two MRI sessions were included. Compared to the synthetic dataset, ADNI exhibits higher inter-subject variability and more pronounced class imbalance, providing a challenging real-world testbed for evaluating the model generalizability.
@@
Since this work focuses on early AD detection, we conduct classification tasks targeting the prodromal stages of impairment: NO vs. MCI and NO vs. SCD. We consider $n_{i}\in\{2,3,4\}$ to reflect subjects with different numbers of visits. Evaluation metrics include accuracy, sensitivity, specificity, precision, F1-score and area under the ROC.
@@
Figure 4: Performance of DiGAN for NO vs. SCD (synthetic) and NO vs. AD (ADNI) across profile lengths of 2, 3, and 4 visits.
@@ Results
Fig. 3 shows the classification performance of DiGAN for distinguishing cognitively normal from subjects with mild cognitive impairment. In the synthetic dataset, DiGAN achieves accuracy above 0.96 and sensitivities exceeding 0.99 for both male and female subjects. While performance remains stable across 2-4 longitudinal visits, a minor decline in specificity is observed for longer sequences in males, potentially due to reduced sample diversity. In contrast, the ADNI dataset, which contains greater inter-subject and acquisition variability, shows a more gradual improvement with increased temporal context: accuracy rises from 0.70 (2 visits) to 0.77 (3 visits), accompanied by similar gains in sensitivity and F1-score. This improvement demonstrates the model’s ability to leverage temporal continuity in longitudinal profiles for improved discrimination of subtle cognitive decline. In addition, Table 2 shows that DiGAN achieves an overall accuracy of 0.948 (synthetic) and 0.710 (ADNI), outperforming all baseline models.
Table 2: Performance comparison of different models on synthetic and ADNI datasets for NO vs. MCI.
@@
Synthetic Accuracy 0.825 $\pm$ 0.055 0.433 $\pm$ 0.024 0.408 $\pm$ 0.044 0.835 $\pm$ 0.043 0.496 $\pm$ 0.031 0.450 $\pm$ 0.028 0.522 $\pm$ 0.064 0.948 $\pm$ 0.016
Sensitivity 0.919 $\pm$ 0.083 0.000 $\pm$ 0.000 0.121 $\pm$ 0.123 0.765 $\pm$ 0.083 0.289 $\pm$ 0.095 0.110 $\pm$ 0.074 0.320 $\pm$ 0.072 0.998 $\pm$ 0.004
Specificity 0.732 $\pm$ 0.036 0.866 $\pm$ 0.027 0.700 $\pm$ 0.121 0.908 $\pm$ 0.024 0.704 $\pm$ 0.059 0.787 $\pm$ 0.039 0.730 $\pm$ 0.103 0.896 $\pm$ 0.031
AUC 0.918 $\pm$ 0.033 0.337 $\pm$ 0.016 0.405 $\pm$ 0.033 0.940 $\pm$ 0.021 0.533 $\pm$ 0.030 0.501 $\pm$ 0.037 0.499 $\pm$ 0.073 0.999 $\pm$ 0.002
Precision 0.773 $\pm$ 0.035 0.000 $\pm$ 0.000 0.213 $\pm$ 0.186 0.891 $\pm$ 0.034 0.486 $\pm$ 0.091 0.303 $\pm$ 0.160 0.551 $\pm$ 0.105 0.907 $\pm$ 0.022
F1-score 0.839 $\pm$ 0.052 0.000 $\pm$ 0.000 0.149 $\pm$ 0.137 0.820 $\pm$ 0.052 0.357 $\pm$ 0.090 0.159 $\pm$ 0.100 0.398 $\pm$ 0.068 0.950 $\pm$ 0.014
@@
NO vs. SCD and NO vs. AD Performance
Fig. 4 shows the classification performance of DiGAN for distinguishing cognitively normal individuals from those with subjective cognitive decline (synthetic) and from Alzheimer’s disease patients (ADNI). Distinguishing NO from SCD remains inherently challenging due to the subtle and self-reported nature of cognitive symptoms. Nevertheless, DiGAN achieves strong discrimination, achieving accuracy above 0.90 and F1-score exceeding 0.92 across most profile lengths. For male subjects, accuracy peaks around three visits with values near 0.94, and female subjects maintain comparable or slightly higher sensitivity, suggesting better generalization across profiles. In the ADNI dataset, where NO is distinguished with clinically diagnosed AD, DiGAN shows clear separability, achieving accuracy above 0.85 and sensitivity exceeding 0.95. As profile length increases from 2 to 4 visits, sensitivity improves consistently, substantiating the model’s ability to exploit longitudinal dynamics. Female subjects show marginally higher sensitivity and F1-scores, indicating that DiGAN effectively captures nuanced temporal biomarkers associated with early neurodegeneration.
@@
Figure 7: ROC and PR curves comparing the DiGAN performance for NO vs. MCI (synthetic).
Fig. 7 presents the ROC and precision–recall (PR) curves comparing DiGAN with baseline methods on the synthetic dataset for NO vs. MCI (male subjects). DiGAN consistently outperforms all state-of-the-art models, with the ROC curve indicating good discrimination between classes even under high specificity constraints. In the PR analysis, DiGAN achieves substantially higher average precision across all recall levels, maintaining stable performance in high-recall regions where baselines degrade rapidly. This robustness indicates that the diffusion-guided representation effectively captures subtle structural and temporal deviations associated with early cognitive decline. Among baselines, GP and TVAE perform moderately well but exhibit trade-offs between precision and recall. DiGAN’s advantage arises from the integration of the diffusion process, which enriches the latent space by reconstructing intermediate neuroanatomical trajectories, and the attention mechanism, which selectively amplifies clinically relevant temporal dependencies.
@@ Conclusion
In this paper, we present DiGAN, which integrates a latent diffusion model with an attention-based convolutional encoder for early detection of AD. By generating realistic longitudinal neuroimaging trajectories from limited training data, DiGAN addresses the scarcity and irregularity of clinical follow-ups and improves temporal continuity in disease modeling. Experiments on synthetic and ADNI datasets show that DiGAN consistently outperforms existing baselines, while providing interpretable embedding and attention maps that align with known neuroanatomical markers of early AD progression.
@@ Acknowledgments
We are thankful to Gabriel Ziegler (DZNE) for providing access to the synthetic dataset and for his valuable feedback throughout this study.

### v2

@@ Abstract
Early diagnosis of Alzheimer’s disease (AD) remains a major challenge due to the subtle and temporally irregular progression of structural brain changes in the prodromal stages. Existing deep learning approaches require large longitudinal datasets and often fail to model the temporal continuity and modality irregularities inherent in real-world clinical data. To address these limitations, we propose the Diffusion-Guided Attention Network (DiGAN), which integrates latent diffusion modelling with attention-guided convolutional network. The diffusion model synthesizes realistic longitudinal neuroimaging trajectories from limited training data, enriching temporal context and improving robustness to unevenly spaced visits. The attention-convolutional layer then captures discriminative structural–temporal patterns that distinguish cognitively normal subjects from those with mild cognitive impairment and subjective cognitive decline. Experiments on the ADNI dataset demonstrate that DiGAN outperforms existing state-of-the-art baselines, showing its potential for early-stage AD detection.
@@ Introduction
  • We demonstrate the effectiveness of DiGAN compared to existing state-of-the-art methods on the ADNI dataset.
@@ 3. Batch Normalization and Output:
A longitudinal study widely used for Alzheimer’s disease research. Subjects with at least two MRI sessions were included. It exhibits higher inter-subject variability and more pronounced class imbalance, providing a challenging real-world testbed for evaluating the model generalizability. Each profile comprises volumetric measures of cortical and subcortical regions, including the entorhinal cortex, parahippocampal gyrus, precuneus, posterior cingulate, and temporal lobes, as well as white matter hyperintensities (WMH) in the frontal and parietal lobes, perivascular spaces (PVS) in the basal ganglia and centrum semiovale.
@@
Since this work focuses on early AD detection, we conduct classification tasks targeting the prodromal stages of impairment: NO vs. MCI and NO vs. AD. We consider $n_{i}\in\{2,3,4\}$ to reflect subjects with different numbers of visits. Evaluation metrics include accuracy, sensitivity, specificity, precision, F1-score and area under the ROC.
@@
Figure 4: Performance of DiGAN for NO vs. AD across profile lengths of 2, 3, and 4 visits.
@@ Results
Fig. 3 shows the classification performance of DiGAN for distinguishing cognitively normal from subjects with mild cognitive impairment. In the ADNI dataset, which contains greater inter-subject and acquisition variability, shows a more gradual improvement with increased temporal context: accuracy rises from 0.70 (2 visits) to 0.77 (3 visits), accompanied by similar gains in sensitivity and F1-score. This improvement demonstrates the model’s ability to leverage temporal continuity in longitudinal profiles for improved discrimination of subtle cognitive decline. In addition, Table 2 shows that DiGAN achieves an overall accuracy of 0.710 (ADNI), outperforming all baseline models.
Table 2: Performance comparison of different models on the ADNI dataset for NO vs. MCI.
@@
NO vs. AD Performance
Fig. 4 shows the classification performance of DiGAN for distinguishing cognitively normal individuals from Alzheimer’s disease patients (ADNI). The DiGAN shows clear separability, achieving accuracy above 0.85 and sensitivity exceeding 0.95. As profile length increases from 2 to 4 visits, sensitivity improves consistently, substantiating the model’s ability to exploit longitudinal dynamics. Female subjects show marginally higher sensitivity and F1-scores, indicating that DiGAN effectively captures nuanced temporal biomarkers associated with early neurodegeneration.
@@
Figure 7: ROC and PR curves comparing the DiGAN performance for NO vs. MCI.
Fig. 7 presents the ROC and precision–recall (PR) curves comparing DiGAN with baseline methods for NO vs. MCI (male subjects). DiGAN consistently outperforms all state-of-the-art models, with the ROC curve indicating good discrimination between classes even under high specificity constraints. In the PR analysis, DiGAN achieves substantially higher average precision across all recall levels, maintaining stable performance in high-recall regions where baselines degrade rapidly. This robustness indicates that the diffusion-guided representation effectively captures subtle structural and temporal deviations associated with early cognitive decline. Among baselines, GP and TVAE perform moderately well but exhibit trade-offs between precision and recall. DiGAN’s advantage arises from the integration of the diffusion process, which enriches the latent space by reconstructing intermediate neuroanatomical trajectories, and the attention mechanism, which selectively amplifies clinically relevant temporal dependencies.
@@ Conclusion
In this paper, we present DiGAN, which integrates a latent diffusion model with an attention-based convolutional encoder for early detection of AD. By generating realistic longitudinal neuroimaging trajectories from limited training data, DiGAN addresses the scarcity and irregularity of clinical follow-ups and improves temporal continuity in disease modeling. Experiments on the ADNI dataset show that DiGAN consistently outperforms existing baselines, while providing interpretable embedding and attention maps that align with known neuroanatomical markers of early AD progression.
@@ Acknowledgments
We are thankful to Gabriel Ziegler (DZNE) and team for his valuable feedback throughout this study.
Data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial magnetic resonance imaging (MRI), positron emission tomography (PET), other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of mild cognitive impairment (MCI) and early Alzheimer’s disease (AD).